### v2

# Dynamic pressure enhancement upon disk impact on a boiling liquid

Abstract
We experimentally investigate the impact of a flat, horizontal disk onto a boiling liquid, i.e., a liquid in thermal equilibrium with its vapor phase. We observe exceptionally high impact pressures deviating strongly from the inertial scaling found for impact in a non-condensable environment, coinciding with the rapid collapse of the vapor pocket entrapped below the disk. We explain our findings, which are relevant for the safe transportation of cryogenic fuels, as a result of vapor condensation, leading to accelerated vapor pocket contraction at high impact velocity and low vapor density.
In any solid-liquid impact, whether one deals with a flat [ermanyuk2011experimental, peters2013splash, jain2021air] or curved [marston2011bubble, hicks2012air, carrat2023air] solid body impacting onto a liquid pool, or with the impact of liquid onto a solid surface [bagnold1939interim, bouwhuis2012maximal, tran2013air, josserand2016drop], air entrapment is a ubiquitous and unavoidable phenomenon, known to provide air-cushioning that mitigates the maximum impact load exerted on the solid body [verhagen1967impact, jain2021air, kim2021water, fan2024air]. It is widely agreed upon that this occurs due to the pressure build-up in the intervening air layer that deforms the liquid surface before impact, resulting in the entrapment of air by the pinch-off of the enclosed air pocket upon impact [lee2012does, bouwhuis2015initial, hendrix2016universal]. While the entrapment of non-condensable air effectively cushions solid-liquid impact, this may no longer hold during boiling liquid impact, where the liquid is in thermal equilibrium with its vapor phase and a condensable vapor pocket is entrapped. With a boiling liquid, a minute increase of pressure –as happens during an impact event– can disturb the equilibrium and may induce a phase change.
Understanding the dynamics of boiling liquid impact is crucial for the safe transportation of cryogenic fuels, such as liquified natural gas (LNG) or liquid hydrogen (LH2) which are pivotal to the energy transition. In addition, this Letter is relevant to the development of LH2-fueled air- and spacecraft, water-hammer phenomena in pipe flow and the fundamental understanding of the role of phase change in multiphase flows. Cryogenic fuels are usually transported in containment tanks at atmospheric or increased pressure, such that they are always in a boiling state, i.e., close to thermodynamic equilibrium with their vapor phase. Slamming of the liquid on the containment wall, commonly known as sloshing, may induce a sudden high impact load during transportation [dias2018slamming]. Experimental studies on boiling liquid impact, i.e. solid-liquid impact in a liquid-vapor system are limited and are mostly at the larger scale of sloshing wave impact [maillard2009influence, lee2021experimental]. Theoretical analysis of the pre-impact stage of LNG-solid impact that assumed a viscous vapor lubrication film suggested that phase change may reduce the pressure buildup in the gas layer, causing less deceleration of the liquid prior to impact [hicks2018lng]. With this, it was postulated that a higher impact pressure may be induced upon impact as less momentum is dissipated before impact.
In this Letter, we experimentally demonstrate that condensation of the entrapped vapor pocket can lead to an increase in excess of an order of magnitude of the local impact pressure on the impacting solid body and, using scaling arguments, in addition explore the conditions under which such an increase may occur. In contrast to [hicks2018lng], we highlight the inertial, condensation-induced collapse of the vapor layer and its effect on enhancing the impact pressure. Here, it is good to note that the reduced Reynolds number ($Re_{\text{lub}}\sim\rho_{\text{v,0}}U_{0}h_{0}/\mu_{\text{v,0}}$) is much greater than 1 in our Letter [jain2021KH], and hence, the vapor flow is inviscid.
To study the dynamics of boiling liquid impact, we build an experimental setup consisting of a 30 $\times$ 30 $\times$ 30 cm sealed chamber with circulating channels on its walls as depicted in Fig. 1a. This allows the control of the ambient temperature $T_{0}$ within the chamber by circulating water with a chiller. To achieve the boiling state, in which the liquid is in thermal equilibrium with merely its own vapor, the chamber is first evacuated with a vacuum pump and flushed with nitrogen gas $(\text{N}_{2})$ twice to eliminate as much moist air as possible from the chamber. Then, the working fluid, Novec 7000 (C4F7OH3) is released into the evacuated chamber from a sealed reservoir. Novec 7000 is used owing to its low saturation vapor pressure at room temperature and its low boiling point. Its properties have been characterized to a sufficient extent in the literature [3MNovec, widiatmo2001equations, ohta2001liquid, rausch2015density, perkins2022measurement, aminian2022ideal] and are summarized in the Supplemental Material [supplemental]. Liquid Novec 7000 vaporizes immediately upon entering the evacuated chamber and the pressure in the chamber $p_{\text{v,0}}$ increases towards its saturation pressure at the desired ambient temperature $T_{0}$. Subsequently, liquid Novec 7000 will start to fill the chamber.
[…]
A circular horizontal flat disk of radius $R_{0}=$ 40 mm and made of steel (RVS 316) is mounted to a linear motor that controls the impact velocity $U_{0}$ of the disk through a rod and is impacted vertically upon the liquid bath with its impacting surface parallel to the free liquid surface. The impact velocity $U_{0}$ is varied from 0.5 m/s to 2.0 m/s at 0.25 m/s intervals. In addition, five different ambient temperatures ranging from $T_{0}$ $\approx$ 11.5 - 24.5 C, are tested in the experiment with the corresponding saturation pressure ranging from $p_{\text{v,0}}$ $\approx$ 420 - 740 mbar. One thing to note is that the saturation pressure reached in our system is always higher than the reference saturation pressure reported in [widiatmo2001equations] by approximately 50 mbar, which we believe is largely due to air dissolved in the liquid. However, due to the large density difference, air is expected to be stabilized above the Novec 7000 vapor and has minimal influence on the vapor entrapment at the liquid-gas interface.
A Kistler pressure sensor (Type 601C) is flush mounted at the disk center to measure the impact pressure at an acquisition rate of 200 kHz. By illuminating light through a diffusive panel, we visualize the free surface below the disk with the total internal reflection technique. The refractive index of Novec 7000 is measured to be 1.25 using the method described in [gluck2011simple] and a minimum liquid height $H=$ 16.5 cm is required to achieve total internal reflection (TIR) for our setup configuration. The reflected free surface is recorded with a high-speed camera at 30k or 48k fps. We synchronize the pressure signal with the high-speed image recording using the procedure outlined in the Supplemental Material [supplemental], with an accuracy of twice the interframe time of the camera.
In Fig. 1b, we plot the time evolution of the pressure $P_{\text{c}}$ measured at the disk center for the smallest ambient temperature reached, $T_{0}\approx 11.5^{\circ}$C, for two different impact velocities, namely $U_{0}=0.5$ m/s (red curve) and $U_{0}=2.0$ m/s (blue). Upon impact at $t_{\text{a}}=0$, both central pressure signals start to build up prominently, and the time delay is consistent with the impact velocity difference. Subsequently, there is a huge difference in the magnitude, where the maximum pressure reached for $U_{0}=0.5$ m/s, $P_{\text{c,max}}=0.084$ bar, is dwarfed by that measured for $U_{0}=2.0$ m/s, namely $P_{\text{c,max}}=19.0$ bar, i.e., more than two orders of magnitude larger.
To understand the origin of this huge difference, Fig. 1c shows a series of representative (re-aspected) snapshots that highlight the distinct behaviour of the entrapped vapor pocket under these different impact conditions. The snapshots are re-aspected to a circular shape, since in the reflected images obtained with the TIR optical set-up that view the free surface at an angle, the disk is deformed into an ellipse. The time after impact $t_{\text{a}}$ is normalized by the inertial time scale $t_{\text{i}}=R_{0}/U_{0}$ to obtain the non-dimensionalized time $\tilde{t}$. In Fig. 1c, the dark region corresponds to the diffuse reflection of the wetted disk (solid-liquid contact) and the inner bright region within the disk perimeter is the mirror reflection of the liquid surface of the entrapped vapor pocket (solid-vapor contact). From the bottom view images, it is evident that a vapor layer is entrapped under the disk after initial impact at $\tilde{t}=0.0063$ at both low ($U_{0}=0.5$ m/s) and high ($U_{0}=$ 2.0 m/s) impact velocity. The mechanism by which the vapor layer is entrapped is the same as that for air layer entrapment, where local pressure build-up in the vertically decelerating intervening gas layer deforms the liquid surface when the solid object is approaching [bouwhuis2015initial, jain2021air]. At low-impact velocity, after impact the liquid-vapor contact line retracts toward the disk center gradually and stably from $\tilde{t}=0.0063$ to $\tilde{t}=0.03$ as seen in Fig. 1c(i) and the pocket is punctured at the center only at a very late stage after impact ($t_{\text{a}}>$ 10 ms as seen in the Supplementary video 1 [supplemental]). In contrast, at high impact velocity, whereas at first the liquid-vapor contact line also starts to retract upon impact, shortly thereafter, (at $\tilde{t}=0.0125$), the entrapped vapor pocket starts to violently collapse from the center region of the disk, resulting in the dark shaded inner region that expands rapidly with time as shown in Fig. 1c(ii). Estimating the spreading speed $U_{\text{sp}}$ of the vapor pocket from the time evolution of the wetted area, we arrive at $U_{\text{sp}}\approx 600$ m/s, which exceeds the speed of sound in liquid Novec 7000, which lies around $C_{\text{sound,L}}\approx 500$ m/s [aminian2022ideal].
Figure 2: Maximum central impact pressure $P_{\text{c,max}}$, rescaled with the inertial pressure scale $\rho_{\text{L}}U_{0}^{2}$ and plotted against the impact velocity $U_{0}$ at different ambient temperature $T_{0}$. Error bars show the standard deviation of the maximum impact pressure at the disk center over at least 5 repetitions of the experiment. The horizontal dashed blue line represents a fit to the data for water–air disk impact from [jain2021air] that follows the classic $\rho_{\text{L}}U_{0}^{2}$ scaling.
In general, the maximum local impact pressure $P_{\text{c,max}}$ at the disk center originates from the liquid stagnation pressure and is therefore expected to follow the classic inertial scaling $\rho_{\text{L}}U_{0}^{2}$, which is confirmed experimentally for water-air disk impact [jain2021air]. In Fig. 2 the rescaled maximum local impact pressure at the disk center, $P_{\text{c,max}}/(\rho_{\text{L}}U_{0}^{2})$, is plotted versus $U_{0}$ for the boiling liquid case at different values of the ambient temperature $T_{0}$. At high ambient temperatures $\left(T_{0}\approx 21.5-24.5^{\circ}\text{C}\right)$, also here the data closely follows the $\rho_{\text{L}}U_{0}^{2}$ scaling, evidenced by data lying on a horizontal line in Fig. 2. Here, the maximum central impact pressures for these higher $T_{0}$ cases are below the horizontal line due to the larger density difference between Novec 7000 vapour and liquid as compared to that of water and air, resulting in a thicker entrapped pocket under the disk which gives more effective cushioning [fan2025disk]. However, as $T_{0}$ decreases, we observe a significant pressure increase, causing $P_{\text{c,max}}$ to strongly deviate from the classic $\rho_{\text{L}}U_{0}^{2}$ scaling law. These exceptionally high impact pressures are always recorded when the vapor pocket was observed to collapse rapidly upon impact from the bottom view. Moreover, the deviation is initiated at a lower impact velocity as the ambient temperature decreases. For instance, at $T_{0}\approx 16.5^{\circ}\text{C}$, pressures start to deviate at $U_{0}=$ 1.5 m/s while $P_{\text{c,max}}$ rises noticeably from $U_{0}=$ 1.0 m/s onwards at $T_{0}\approx 11.5^{\circ}\text{C}$.
Figure 3: Schematic of the model for the collapse of the entrapped vapor pocket upon impact. As the disk descends at constant impact velocity $U_{0}$ the pressure $p_{\text{v}}$ and temperature $T_{\text{v}}$ within the vapor pocket increase, whereas the temperature of the liquid bulk remains at the saturation level ($T_{0}$). Therefore, to restore equilibrium, the vapor pocket starts to condense at the liquid-vapor interface where the released latent heat is conducted into the liquid. Eventually, this process may cause the vapor pocket to collapse and the wetted area spreads in the radial direction over the disk surface.
We attribute the coinciding strong increase of the impact pressure and the violent collapse of the vapor pocket to a rapid condensation of the vapor inside the pocket, which is a distinguishing characteristic of vapor from air [plesset1977bubble, prosperetti2017vapor]. To rationalize this effect, we assume that upon impact a thin vapor pocket of thickness $h_{0}$ has been entrapped below the disk, which subsequently becomes pressurized by the liquid, as sketched in Fig. 3. In response to pressurization, the vapor in the pocket will condense on the liquid-vapor interface, where latent heat is produced that needs to be transported into the liquid [plesset1977bubble]. As further justified in the Supplemental Material [supplemental], neglecting (i) sensible heat in the vapor, (ii) heat transport through the vapor phase, (iii) density changes in the liquid phase and (iv) condensation on the disk we may write
$L\frac{dm_{\text{v}}}{dt}=-S_{\text{A}}k_{\text{L}}\left.\frac{\partial T}{\partial n}\right|_{\text{interface}}\approx S_{\text{A}}k_{\text{L}}\frac{\Delta T_{\text{v}}}{\delta_{\text{th}}}\,$, (1)
where $L$ is the latent heat of vaporization, $dm_{\text{v}}/dt$ the rate of change of the vapor mass in the pocket, $S_{\text{A}}$ the surface area of the pocket, and $k_{\text{L}}=\alpha_{\text{L}}\rho_{\text{L}}C_{\text{L}}$ the thermal conductivity of the liquid (with $\alpha_{\text{L}}$ and $C_{\text{L}}$ the thermal diffusivity and specific heat capacity of the liquid, respectively). The normal temperature gradient $\partial T/\partial n$ at the interface can be estimated as the ratio of the temperature difference $\Delta T_{\text{v}}=T_{\text{v}}-T_{0}$ between vapor and liquid and the thermal boundary layer thickness $\delta_{\text{th}}\approx\sqrt{\pi\alpha_{\text{L}}t_{\text{c}}}$ where $t_{\text{c}}$ is the collapse time scale of the vapor pocket defined as $t_{\text{c}}\approx h_{0}/U_{0}$. The right hand side of Eq. (1) can be rewritten as
$\frac{dm_{\text{v}}}{dt}\approx-\rho_{\text{v,0}}S_{\text{A}}U_{\text{cond}}\,$, (2)
where $\rho_{\text{v,0}}$ is the equilibrium vapor density and $U_{\text{cond}}$ is the vapor flux into the liquid interface, i.e., the speed at which vapor condenses on the pocket wall.
The temperature rise $\Delta T_{\text{v}}$ in the vapor phase can be estimated from the pressure increase $\Delta p_{\text{v}}$ upon pressurization using the linearized Clausius-Clapeyron equation
$\frac{\Delta T_{\text{v}}}{T_{0}}\approx\beta\frac{\Delta p_{\text{v}}}{p_{\text{v,0}}}\,$, (3)
where $\beta=R_{\text{s}}T_{0}/L$, with $R_{\text{s}}$ the specific gas constant of the vapor, is a small dimensionless parameter. Combining Eqs. (1),  (2), and (3), the vapor flux into the liquid interface can be written as
$\frac{U_{\text{cond}}}{U_{0}}\sim\frac{\beta^{2}}{\sqrt{\pi}}\frac{\rho_{\text{L}}}{\rho_{\text{v,0}}}\frac{C_{\text{L}}}{R_{\text{s}}}\sqrt{\frac{\alpha_{\text{L}}}{h_{0}U_{0}}}\frac{\Delta p_{\text{v}}}{p_{\text{v,0}}}\,.$ (4)
Figure 4: Plotting the rescaled maximum impact pressure $P_{\text{c,max}}/\rho_{\text{L}}U_{0}^{2}$ at the disk center from Fig. 2 against $U_{\text{sp}}/C_{\text{sound,L}}$ as derived from the proposed theoretical model gives a decent collapse of data. The maximum central impact pressure induced during boiling impact can be described as a result of vapor condensation that strongly depends on the impact velocity and vapor density of the boiling system. (Note that an additional data point for $U_{0}=2.25$ m/s at $T_{0}\approx 14.5^{\circ}$C is included here.)
[…]
Finally, while condensation occurs along the entire vapor pocket interface, the collapse of the entrapped vapor pocket is determined by how fast the liquid-solid contact area spreads in the plane of the disk surface. Using continuity to relate the spreading speed of the liquid-solid contact area $U_{\text{sp}}$ over the disk surface to $U_{\text{cond}}$, namely $2U_{\text{sp}}h_{0}\sim U_{\text{cond}}R_{0}$, we obtain
$\frac{U_{\text{sp}}}{C_{\text{sound,L}}}\sim\frac{\beta^{2}}{2\sqrt{\pi}}\left(\frac{R_{0}}{h_{0}}\right)^{\!\!3/2}\!\!\!\frac{\rho_{\text{L}}}{\rho_{\text{v,0}}}\frac{C_{\text{L}}}{R_{\text{s}}}\sqrt{\frac{\alpha_{\text{L}}}{R_{0}U_{0}}}\frac{\rho_{\text{L}}U_{0}^{2}}{p_{\text{v,0}}}\,.$ (5)
In this equation, the only remaining unknown quantity is the initial thickness $h_{0}$ of the vapor pocket. Prior to impact, when the disk is moving towards the free liquid surface, pressure builds up as $\sim\rho_{\text{v}}U_{0}^{2}$ underneath the impacting disk as the intervening vapor layer between the disk and the liquid surface is being squeezed. This causes the liquid surface to deform, with the surface under the disk center being depressed the most [mayer2018flat, jain2021KH] and the surface near the disk edge being lifted due to Kelvin-Helmholtz instability [jain2021KH]. Therefore, as illustrated in Fig. 3, upon impact, the disk will first contact with the deformed liquid surface near the disk edge and entrap a thin layer of vapor with an initial radius $R_{0}$ and central thickness $h_{0}$ of the order of hundreds of microns. In the potential flow framework [peters2013splash, bouwhuis2015initial], one expects the deformation of the liquid interface to be proportional to the vapor to liquid density ratio, such that $h_{0}/R_{0}=\lambda\rho_{\text{v,0}}/\rho_{\text{L}}$, where $\lambda$ is a numerical constant of order unity. Inserting this expression in Eq. (5) gives
$\frac{U_{\text{sp}}}{C_{\text{sound,L}}}\sim\frac{\beta^{2}}{2\sqrt{\pi}\lambda^{3/2}}\left(\frac{\rho_{\text{L}}}{\rho_{\text{v,0}}}\right)^{\!\!5/2}\!\!\frac{C_{\text{L}}}{R_{\text{s}}}\sqrt{\frac{\alpha_{\text{L}}}{R_{0}U_{0}}}\frac{\rho_{\text{L}}U_{0}^{2}}{p_{\text{v,0}}}\,.$ (6)
In the above expression the dependence on the impact velocity $U_{0}$ is clear since $U_{\text{sp}}\sim U_{0}^{3/2}$. Lowering the ambient temperature $T_{0}$, the dominant effect is in the reduced vapor density $\rho_{\text{v,0}}$, which influences $U_{\text{sp}}$ both directly and through a smaller initial thickness $h_{0}$ of the entrapped vapor layer.
We then compute $U_{\text{sp}}/C_{\text{sound,L}}$ using values of the temperature-dependent properties in Eq. (6) from [aminian2022ideal, widiatmo2001equations, rausch2015density], see Supplementary Material [supplemental]. The fitting parameter $\lambda$ is taken to be $2.1$ as obtained by fitting the experimentally measured $h_{0}$ (not shown here, but see [fan2025disk]). Replotting the rescaled maximum impact pressure at the disk center from Fig. 2 against the computed $U_{\text{sp}}/C_{\text{sound,L}}$ gives a decent collapse of data as shown in Fig. 4, suggesting that this simple model predicts the dependence of the vapor condensation in the pocket on the impact velocity and ambient temperature during boiling liquid impact. In addition, this confirms that the condensation-induced collapse of the vapor pocket is responsible for the very high impact pressures exerted on the impacting surface.
The mechanism by which condensation creates the anomalously high maximum central impact pressures is the following: When phase change does not play a dominant role, the solid-liquid impact is cushioned by the (relatively slow) pressurization of the vapor pocket. If the vapor condenses sufficiently fast, the liquid reaches the solid interface unbraked, resulting in a large and sudden impact pressure peak.
In summary, we experimentally observed anomalously high impact pressures during impact of a disk onto a boiling liquid, exceeding expected values by more than an order of magnitude. These high pressures are observed to coincide with a rapid collapse of the vapor pocket entrapped below the disk, and found to occur for low equilibrium vapor densities and high impact velocities. We constructed a scaling model accounting for our findings that provides quantitative insight in the material parameter values for which anomalous behaviour may occur. Most notably, our experiments show that phase change occurring during impact under thermal equilibrium conditions may lead to qualitatively different physics compared to that encountered during impact experiments in non-condensable gases such as air. Clearly, vapor condensation during boiling liquid impact is detrimental to the effectiveness of gas cushioning and may lead to unexpectedly high local impact pressure on the solid surface upon impact. Therefore, it is imperative to account for phase change during boiling liquid impact while designing a containment system for cryogenic fuels to ensure the safety and integrity of the tank.
We thank G.-W. Bruggert, D. van Gils, M. Bos and T. Zijlstra for their technical support. We would also like to thank U. Jain for providing his original experimental data. This work is part of the Vici project IMBOL (project number 17070) which is partly financed by the Dutch Research Council (NWO).
Data availability: The data that support the findings of this article are openly available [dataSet].